OpenStudy (anonymous):

Greg made two similar glass display boxes shaped like triangular prisms. The length of the sides of the base of the first display box is 5 feet, and the length of the sides of the base of the second display box is 10 feet. If the volume of the first display box is 52 ft3, what is the volume of the second display box?

OpenStudy (anonymous):

104 i think

OpenStudy (anonymous):

5/52:10/x 5x=52(10) 5x=520 x=520/5 x=104

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

you see while calculating the volume of prism.. it is (base area) * height But base area = area of triangle = sqrt(3) /4 * a^2... where a is the side of the triangular base So, when the side is doubled.. the volume and the base area is quadrupled. So the correct answer will be 52*4= 208 cubic feet

OpenStudy (anonymous):

ty, and your sure that is the correct answer because both are choices

OpenStudy (anonymous):

yes.. I am.. don't worry about that..

OpenStudy (anonymous):

first one is for those who don't pay enough attention and fall into the trap the question has set.. believe me.. I have taught high school students for some years to know this..